consider a mass of m1 and m2. m2 larger than m1. note that gravity acts on both masses. determine the net force acting on the system.
what is the total mass of the system?
with the info you just computed fine the accerleration of the system.
1
Expert's answer
2011-10-25T10:55:07-0400
Consider a mass of m1 and m2. m2 larger than m1. Note that gravity acts on both masses.

1). Determine the net force acting on the system.

The force acting on m1 is g*m1 and the force acting on m2 is g*m2. The net force acting on the system is F = g*m1+g*m2 = g*(m1+m2).

2). What is the total mass of the system?

The total mass of the system is M = m1+m2.

3). With the info you just computed find the acceleration of the system.

According to Newton's second law, a = F/M = g*(m1+m2)/(m1+m2) = g.